# VSC-RL: Advancing Autonomous Vision-Language Agents with Variational Subgoal-Conditioned Reinforcement Learning

### v1

@@ Appendix A Implementation Details
As shown in Table 2, we summarize VLMs used in VSC-RL. For MiniGrid, we built our VSC-RL on the open repository of babyAI (Chevalier-Boisvert et al., 2019), hyper-parameter settings are listed in Table 3. For AitW tasks, we built our VSC-RL on the open repository of DigiRL (Bai et al., 2024), hyper-parameter settings are listed in Table 4.
Table 2: The summary of VLMs used in VSC-RL.
Component Name VLM Type Description
Subgoal Generator VLM Gemini-1.5-Pro (Team et al., 2024) Autonomously decompose the goal to the subgoals.
Reference Actor $\pi_{\text{ref}}$ AutoUI-Base (Zhang & Zhang, 2023) Provide reference action for imitation.
Target Actor $\pi$ AutoUI-Base Make decisions to maximize subgoal-conditioned return.
Evaluator (Pan et al., 2024b) Gemini-1.5-Pro Autonomously evaluate the goal’s or subgoal’s success.
Table 3: Hyper-parameters setting of VSC-RL on MiniGrid.
Hyper-parameter Value
Batch Size 256
Total Steps 200,000
Discount Factor 0.99
Learning Rate 1e-3
Network Layers (Image) 3
Network Layers (Text) 1
Network Layers (Actor) 2
Network Layers (Critic) 2
Activation ReLU
Optimizer Adam
Table 4: Hyper-parameters setting of VSC-RL on AitW tasks.
Hyper-parameter Value
@@
Update Epoch (Actor) 20
@@ Appendix B Theoretical Analysis
$\mathop{\mathbb{E}}_{\tau\sim p_{\pi}(\tau|g)}\left[\log p(O|\tau,g)\right]\Rightarrow\sum_{i=1}^{N}\left[\mathop{\mathbb{E}}_{\tau_{i}\sim p_{\pi}(\tau_{i}|g_{i})}\left[\log p(O|\tau_{i},sg_{i})\right]\right]$
@@
Given goal-conditioned reference policy $\pi_{\text{ref}}$ and subgoal-conditioned target policy $\pi$, the goal-conditioned KL divergence of a given goal $g$ has the upper bound of subgoal-conditioned KL divergence of corresponding subgoals $\{sg_{i}\}_{i=1}^{N}$ as followed:
@@
Appendix C Prompt Example
We provide the prompt example of the subgoal generator in our VSC-RL for a given goal and corresponding decomposed subgoals in Figure 11 and Figure 12.
@@ Appendix C Prompt Example
Appendix D Qualitative Example
We provide qualitative examples of VSC-RL applied to MultiRoom (Figure 13), AitW General (Figure 14), and Web Shopping tasks (Figure 15).
@@ Appendix D Qualitative Example
Figure 14: Qualitative example of VSC-RL on the General task.
Figure 15: Qualitative example of VSC-RL on the Web Shopping task.

### v2

@@ Appendix A Implementation Details
As shown in Table 3, we summarize LLMs and VLMs used in VSC-RL. For AitW tasks[27], we built our VSC-RL on the open repository of DigiRL [4]. Hyperparameter settings are listed in Table 4, and each run of VSC-RL takes approximately 24 hours on 1 NVIDIA GeForce RTX 4090 GPU and 8 Intel Xeon CPUs. For WebArena-Lite tasks [17], we built our VSC-RL on the open repository of WebRL [26] and follow the online RL loop of interaction, filtering, and update. To ensure a fair comparison, we remove WebRL’s curriculum learning component, so that all methods are trained solely on the same task set. For VSC-RL and other baselines, we apply the same actor perplexity-based filtering strategy as WebRL to select replay data, ensuring consistency in experience quality. Hyperparameter settings are listed in Table 5, and each run of VSC-RL takes approximately 24 hours on 8 NVIDIA GeForce RTX 4090 GPUs and 8 Intel Xeon CPUs. Specifically, we report some results on AitW (Set-of-Marks, AppAgent, CogAgent, AutoUI, and Filtered BC) and WebArena-Lite (SFT, Filtered BC, and AWR) from the literature [4, 26].
Table 3: The summary of LLMs and VLMs used in VSC-RL.
Component Task Description
AitW WebArena-Lite
Subgoal Generator Gemini-1.5-Pro Gemini-1.5-Pro Autonomously decompose the goal into subgoals.
Reference Actor $\pi_{\text{ref}}$ AutoUI-Base Llama-3.1-8B Provide reference action for imitation.
Target Actor $\pi$ AutoUI-Base Llama-3.1-8B Make decisions to maximize subgoal-conditioned return.
Evaluator Gemini-1.5-Pro Llama-3.1-8B Autonomously evaluate the goal’s or subgoal’s success.
Table 4: Hyperparameters settings of VSC-RL on AitW tasks.
Hyperparameter Value
@@
Table 5: Hyperparameter settings of VSC-RL on WebArena-Lite tasks.
Hyperparameter Value
Batch Size 128
Total Trajectories 1,000
Discount Factor 0.9
Learning Rate 1e-6
Update Epoch (Actor Equation 5) 1
Update Epoch (Critic Equation 6) 1
Update Epoch (Actor Equation 7) 1
Maximum Gradient Norm 1.0
@@ Appendix B Theoretical Analysis
$\max_{\pi}\mathop{\mathbb{E}}_{\tau\sim p_{\pi}(\tau|g)}\left[\log p(O|\tau,g)\right]\Rightarrow\max_{\pi}\sum_{i=1}^{N}\left[\mathop{\mathbb{E}}_{\tau_{i}\sim p_{\pi}(\tau_{i}|g_{i})}\left[\log p(O|\tau_{i},sg_{i})\right]\right]$
@@
Given goal-conditioned reference policy $\pi_{\text{ref}}$ and subgoal-conditioned target policy $\pi$, the goal-conditioned KL divergence of a given goal $g$ has the upper bound of subgoal-conditioned KL divergence of corresponding subgoals $\{sg_{i}\}_{i=1}^{N}$ as follows:
@@
Appendix C Additional Experiments: MiniGrid
We also evaluate our VSC-RL on the toy vision-language decision-making tasks, MiniGrid [9]. We select the PPO [29] as the baseline, and we apply VSC-RL in the PPO for a fair comparison. We built our VSC-RL on the open repository of babyAI [9], hyperparameter settings are listed in Table 6. Overall, as shown in Figure 8, our VSC-RL outperforms the baseline in all tasks remarkably, especially in the difficult task with the increasing number of rooms. From the result of MultiRoom-N2-v0 shown in LABEL:fig:multiroom_2, we can tell that although PPO and VSC-RL both successfully reach $100\%$ success rate, our VSC-RL shows a better sample efficiency. For MultiRoom-N4-v0 (LABEL:fig:multiroom_4) and MultiRoom-N6-v0 (LABEL:fig:multiroom_6) where PPO is not able to learn any useful policy, while VSC-RL exhibits strong performance of $100\%$ and $80\%$ success rate, respectively.
Table 6: Hyperparameter settings of VSC-RL on MiniGrid.
Hyperparameter Value
Batch Size 256
Total Steps 200,000
Discount Factor 0.99
Learning Rate 1e-3
Network Layers (Image) 3
Network Layers (Text) 1
Network Layers (Actor) 2
Network Layers (Critic) 2
Update Epoch (Actor Equation 5) 4
Update Epoch (Critic Equation 6) 4
Update Epoch (Actor Equation 7) 4
Activation ReLU
Optimizer Adam
Appendix D Additional Experiments: Subgoal Generator in VSC-RL
Improvement from Subgoal Generator.
We investigate the importance of the subgoal generator in our VSC-RL on the Web Shopping subsets with different horizon lengths (short, medium and long). We implement VSC-RL with the original goal instead of the subgoals generated from VLM. As shown in Figure 11, the subgoal generator can effectively improve the performance across all types of Web Shopping tasks via autonomously decomposing the original goal into subgoals. Especially, the subgoal generator can effectively enhance the $50\%$ and $32\%$ performance in the Web Shopping medium and long tasks, respectively.
Verification of Subgoal Generator.
To investigate the quality and feasibility of the generated subgoals, we manually verify the results of the subgoal generator on $135$ trajectories from the AitW human demonstration. There are $135(100\%)$ goals that are decomposed into feasible subgoals successfully, and the final goal can be accomplished by reaching these subgoals sequentially. Specifically, there are $123(91.1\%)$ goals that are decomposed into subgoals completely aligning with the human demonstration. For the remaining $12(8.9\%)$ goals, the subgoal generator provides alternative subgoals different from human demonstration, but still can successfully arrive at the final goal.
Appendix E Additional Experiments: WebArena-Lite
As shown in Table 7, we present the evaluation performance on specific WebArena-Lite tasks, including Reddit, Gitlab, CMS (online store content management system), Map (OpenStreetMap), and OSS (OneStopShop). We also provide the learning curves in Figure 17. These empirical results demonstrate that our VSC-RL achieve the best performance across on all types of tasks, significantly surpassing existing SOTAs.
Table 7: The evaluated performance on WebArena-Lite. The best performance is highlighted.
Method Task (# Ratio)
Reddit (12.7%) Gitlab (19.4%) CMS (21.2%) Map (18.8%) OSS (27.9%) All (100.0%)
SFT 36.8% 6.7% 20.0% 33.3% 17.8% 20.6%
Filtered BC 52.6% 20.0% 31.4% 23.3% 8.9% 23.0%
AWR 57.9% 26.7% 31.4% 26.7% 17.8% 28.5%
DigiRL 52.4% 28.1% 37.1% 32.3% 15.2% 30.3%
WebRL 57.1% 28.1% 34.3% 35.5% 15.2% 30.9%
VSC-RL (ours) 61.9% 31.3% 40.0% 35.5% 19.6% 34.5%
Figure 17: Learning curves on (a) Reddit, (b) Gitlab, (c) CMS, (d) Map, (e) OSS, and (f) All tasks.
Appendix F Prompt Example
We provide the prompt example of the subgoal generator in our VSC-RL for a given goal and corresponding decomposed subgoals in the MultiRoom (Figure 18), AitW (Figure 19) and WebArena-Lite (Figure 20) tasks.
@@ Appendix F Prompt Example
Figure 20: Prompt example for our subgoal generator for WebArena-Lite tasks. The subgoal generator decomposes a user task into a sequence of more specific and actionable instructions based on the provided HTML context.
Appendix G Qualitative Example
We provide qualitative examples of VSC-RL applied to MultiRoom (Figure 21), AitW General (Figure 22), AitW Web Shopping (Figure 23), WebArena-Lite Map (Figure 24), and WebArena-Lite CMS (Figure 25).
@@ Appendix G Qualitative Example
Figure 22: Qualitative example of VSC-RL on the AitW General task.
Figure 23: Qualitative example of VSC-RL on the AitW Web Shopping task.
Figure 24: Qualitative example of VSC-RL on the WebArena-Lite Map task.
Figure 25: Qualitative example of VSC-RL on the WebArena-Lite CMS task.